OpenStudy (mikeymaximum):

Solve the equation. x + 1/6 = 6 A. x = –6 1/6 B. x =–5 5/6 C. x = 5 5/6 D. x = 6 1/6

OpenStudy (yolo_king):

x=\[\frac{ 35 }{ 6 }\]

OpenStudy (secret-ninja):

To turn an improper fraction into a mixed fraction you just need to divided the numerator by the denominator. 35 / 6 = 5.83 So you have: \[5 \frac{ ? }{ 6 }\] Can you guess the numerator?

OpenStudy (secret-ninja):

C. 5 5/6 :)